OpenStudy (anonymous):

If f(x)=square root 6sin(x)+9 then f'(0)=

OpenStudy (anonymous):

\[\sqrt{6\sin(x)+9}\]

OpenStudy (anonymous):

\[f(x)=(6\sin{x}+9)^{\frac{ 1 }{ 2 }}\]\[f'(x) = \frac{ 1 }{ 2 }(6\sin{x}+9)^{\frac{ -1 }{ 2 }} \times 6 \cos {x}\]Hence f'(0) =1

OpenStudy (anonymous):

so then the answer is 1 correct?

OpenStudy (anonymous):

@Anonymous1921 is that the correct answer that I posted just making sure

OpenStudy (anonymous):

1 is the answer right?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

ok thanks